Abel and Baker decided to form a partnership. Abel contributed equipment (book value $65,000), inventory (paid $20,000), and $10,000 cash. The equipment and inventory have a current market value of $40,000 and $15,000, respectively. Abel also had a debt of $20,000 for the equipment. Baker contributed office equipment (book value $20,000) and cash of $50,000. The current market value of the office equipment is $10,000. The two partners fail to agree on a profit-and-loss-sharing ratio. For the first month (June), the partnership lost $4,000.

  1. How much of this loss goes to Abel? How much goes to Baker?

  2. The partners withdrew no assets during June. What is each partner’s Capital balance at June 30? Prepare a T-account for each partner’s Capital.

3 Partners’ profits, losses, and Capital balances

2. Baker Capital, $58,000
Answer #1

Solution :

1. If the profit sharing ratio is not decided among the partners , it will be shared equaly between the partners. Hence, Abel would get $ 2000 loss share and Baker would get a loss share of $ 2000

Partners Capital A/c .

Dr. Cr.

Particulars Abel Baker Particulars Abel Baker
To profit and loss appropriation A/c 2000 2000 By Capital 45000* 60000**
To Balance C/d 43000 58000
Total 45000 60000 Total 45000 60000

* Capital of Abel = Market value of equipements + Market value of inventory + Cash - Debt brought by partner against equipement = 40000+ 15000+ 10000- 20000 = 45000

** Capital of baker = Market value of equipment + Cash = 10000 + 50000 = 60000

Balances as on 30 june :

Abel = $ 43000

Baker = $ 58000